# The life cycle of giant molecular clouds in simulated Milky Way-mass galaxies

### v1

@@ 3.6 Star formation efficiency per free-fall time in GMC life cycles
To further follow the SFE in different stages within the life cycles of GMCs, we calculate the star formation efficiency per free-fall time $\varepsilon_{\mathrm{ff}}$ by equation (3) for all tracked GMCs in the SFE1, Rad, SN, and Nofeed runs. We then select all $t_{\mathrm{life}}=10\mathrm{Myr}$ GMCs among them and for each model variation, we divide $t_{\mathrm{life}}=10\mathrm{Myr}$ GMCs into $15$ equal-cloud number bins according to the maximum baryonic mass $M_{\mathrm{max,baryon}}$. Within each maximum baryonic mass bin, we average $\varepsilon_{\mathrm{ff}}$ among GMCs at every moment and get $\varepsilon_{\mathrm{ff}}$ as a function of time $t$ in Fig. 9. Here we note that it is almost impossible to choose a fair value of GMC lifetime in this comparison of $4$ simulations with various feedback setups. On the one hand, we want to have sufficient clouds to be statistically meaningful, so we can’t choose a too large value for a lifetime. On the other hand, we do not want to choose a too short value to miss any physical effect in the late life of clouds such as supernovae and long-term dynamics. Here we choose $10$ Myr as a balance between these limits. In general, GMCs among all of Rad, SN, and Nofeed runs show an increasing $\varepsilon_{\mathrm{ff}}$ in the early stages and decreasing $\varepsilon_{\mathrm{ff}}$ in the last several $\mathrm{Myr}$ of their life cycles. Such a trend is consistent with the results in Grudić et al. (2019). Among these four model variation, GMCs with higher $M_{\mathrm{max,baryon}}$ exhibit higher $\varepsilon_{\mathrm{ff}}$. The high $\varepsilon_{\mathrm{ff}}$ (larger than $2$ percent) in more massive GMCs can sustain for the entire life cycles of these GMCs, while less massive GMCs can have extremely low $\varepsilon_{\mathrm{ff}}$ (less than $2$ percent) in the first few $\mathrm{Myr}$. Compared with both the SFE1 run and the Rad run, the star formation in the SN run is slightly more active across all mass ranges. Given that the selected timescale is $10$ Myrs here, it is relatively short for the supernova to take effects thus the star formation in the SN run is mainly regulated by the galactic environment.
@@ 4.1 Comparisons to previous work
During the past several years, thanks to the increasing power of supercomputers and a better understanding of the ISM physics in galaxy formation simulations, a few studies have investigated the GMC properties from large-scale galaxy simulations. Here we would like to compare our results with them. Regarding the physical properties of identified GMCs, such as mass, virial parameter, effective radius, velocity dispersion, and Larson’s relations, GMCs identified in the Feedback ($3$D) case from Grisdale et al. (2018) in general exhibit similar distributions as shown in our fiducial run. The mass distribution of GMCs in the m12m case of Benincasa et al. (2020) is also similar to that in our fiducial run. Benincasa et al. (2020) also find GMC lifetimes within a couple of million years, consistent with our fiducial run.
@@
In terms of GMC lifetimes, Benincasa et al. (2020), J21, and our work all find that GMC lifetimes range from a couple of $\mathrm{Myr}$ to several tens of $\mathrm{Myr}$. However, Benincasa et al. (2020) report limited dependence of the lifetime on the GMC mass and J21 proposes a negative correlation between GMC lifetimes and masses, which are contradictory to our findings on the positive correlation between GMC lifetimes and masses. In Benincasa et al. (2020), they use a relatively simple tracking algorithm and only follow strictly gravitationally bound GMCs. This approach may overlook the early stages of the GMC life cycles and miss key accretion phases, which are crucial to establishing the positive correlation between GMC lifetimes and masses. As for J21, though they introduce a similar tree network approach to capture both the temporal evolution of clouds and their interactions like our work, here we highlight one significant difference between J21’s and our GMC-tracking method, which may cause different results on the relation of GMC lifetimes and masses. Using the quasi-Lagrangian nature of arepo, we record additional information about the gas mass flow between nodes, which provides physical guidelines for selecting paths among multiple branches. J21, on the contrary, assumes the same probabilities for branches, regardless of the gas mass flowing through them. In this way, the GMC tracking in J21 may be misled by statistically overweighting minor branches of GMC trees. A massive GMC may go through many splitting or merging processes within its lifetime, and thus produce many minor branches. According to their tracking algorithm, these minor branches shorten the characteristic lifetimes of massive GMCs. It should be mentioned that in the GRIFFIN project, only clouds with the longest lifetime can form the most massive star clusters (Lahén et al. 2020), consistent with our findings here.
@@
Regarding the star formation process, while many studies focus on isolated GMC simulations to explore local SFE, our approach captures the influence of galactic environments self-consistently, while still allowing investigation of local SFE. In Section 3.5, we notice that the relation between the cloud surface density and the SFE is in a similar form as found in the isolated GMC simulations (e.g. Grudić et al. 2018; Li et al. 2019). Moreover, the large variations in the local SFE are observed in all simulation runs, consistent with Grisdale et al. (2019). Besides the fiducial run with the stellar feedback, we utilize the GMCs in the Nofeed run to quantitatively approximate the effect of the galactic shear on the local star formation efficiency with $\Sigma_{\mathrm{shear}}$. The best-fitting value of $\Sigma_{\mathrm{shear}}$ is comparable to the best-fitting value of $\Sigma_{\mathrm{fb}}$, suggesting the non-negligible role that the galactic shear plays in the star formation activity. We would like to point out that in Grudić et al. (2018) and Grudić et al. (2023), while their best-fitting $\Sigma_{\mathrm{fb}}$ is close to our best-fitting value, an extra parameter namely $\epsilon_{\mathrm{max}}$ is artificially introduced to account for any possible additional suppression of star formation other than stellar feedback. This approximation may partly mimic the effect of the $\Sigma_{\mathrm{shear}}$ in our study. Similar to our work, Khullar et al. (2024) also track life cycles of GMCs and study the GMC-scale SFE in galactic scale simulations, they get similar results on the importance of stellar feedback regulating star formation and also the large scatter in local SFE.
@@ 4.2 Limitations and future perspectives
We acknowledge several limitations of this work and outline potential directions for future research. First, while arepo, as a moving-mesh code, has a quasi-Lagrangian nature, it is not purely Lagrangian. The timescale for cell refinement and de-refinement is much shorter than the snapshot output cadence (i.e., $1\mathrm{Myr}$). Consequently, using cell IDs to track mass flow between GMCs across different epochs can lead to inaccuracies due to cell reconstruction during these intervals. To address this, future simulations will incorporate the Monte Carlo tracer particles (Genel et al. 2013) to allow tracking the mass flow of clouds more accurately. Second, though we find that galactic shear has a significant effect on GMC-scale star formation, it is yet unclear how such an effect would vary in different galactic environments. To explore how various galactic environments might have different effects on GMC-scale star formation, a companion study (Deng et al., in preparation) is currently underway to explore the evolution of GMCs in dwarf galaxies, utilizing the solar-mass-resolution ISM model, RIGEL (Deng et al. 2024).
@@ 5 Summary
In this study, we track the temporal evolution of a large sample of GMCs within simulated Milky Way-sized galaxies using the hydrodynamic code arepo under the SMUGGLE galaxy formation framework. The high mass resolution ($\sim 10^{3}M_{\odot}$) and high-cadence output frequency (1 Myr) enable the identification of cloud candidates and the construction of evolutionary trees. These evolutionary trees successfully capture the full GMC life cycle, including formation, gas accretion and collapse, star and cluster formation, and cloud dispersal. To investigate how different physical processes influence GMC evolution and star formation on cloud scales, we explore a set of model variations with different star formation and stellar feedback parameters. Below, we summarize the key findings:
@@
  • GMC lifetimes vary significantly across model variations, ranging from a few to several tens of Myr, suggesting a strong influence from the star formation and stellar feedback subgrid models. Higher $\epsilon_{\rm ff}$ leads to faster gas consumption and cloud dispersal due to earlier and stronger stellar feedback and thus shorter lifetime. In the Nofeed run, GMCs live substantially longer than in other models, with over $10$ surviving beyond $20\,\mathrm{Myr}$. In the SN run, GMCs with lifetimes below $20\,\mathrm{Myr}$ tend to last longer than in the fiducial run, whereas those exceeding $20\,\mathrm{Myr}$ have shorter lifetimes. Conversely, in the Rad run, GMCs with lifetimes below $20\,\mathrm{Myr}$ show distributions similar to the fiducial run, while those above $20\,\mathrm{Myr}$ persist longer due to the absence of SN feedback.
@@
  • We capture the entire life cycles of GMCs in our simulated galaxies, from accretion and star formation to stellar feedback-driven cloud dispersal. We find that the accretion process continuously supplements the gas content throughout most of GMCs’ life cycles, making the cloud mass higher and virial parameter lower in the early stage. When the cloud mass reaches its peak, feedback from young stars disperses the cloud, reduces the cloud mass and inflates the virial parameter. Interestingly, there is a positive correlation between GMC lifetimes and GMC maximum gaseous masses, which stands true regardless of star formation and stellar feedback models/parameters.

### v2

@@ 3.6 Star formation efficiency per free-fall time in GMC life cycles
To further follow the SFE in different stages within the life cycles of GMCs, we calculate the star formation efficiency per free-fall time $\varepsilon_{\mathrm{ff}}$ by equation (7) for all tracked GMCs in the SFE1, Rad, SN, and Nofeed runs. We then select all $t_{\mathrm{life}}=10\,\mathrm{Myr}$ GMCs among them and for each model variation, we divide $t_{\mathrm{life}}=10\,\mathrm{Myr}$ GMCs into $15$ equal-cloud number bins according to the maximum baryonic mass $M_{\mathrm{max,baryon}}$. Within each maximum baryonic mass bin, we average $\varepsilon_{\mathrm{ff}}$ among GMCs at every moment and get $\varepsilon_{\mathrm{ff}}$ as a function of time $t$ in Fig. 9. Here we note that it is almost impossible to choose a fair value of GMC lifetime in this comparison of $4$ simulations with various feedback setups. On the one hand, we want to have sufficient clouds to be statistically meaningful, so we cannot choose a too large value for a lifetime. On the other hand, we do not want to choose a too short value to miss any physical effect in the late life of clouds such as supernovae and long-term dynamics. Here we choose $10$ Myr as a balance between these limits. In general, GMCs among all of Rad, SN, and Nofeed runs show an increasing $\varepsilon_{\mathrm{ff}}$ in the early stages and decreasing $\varepsilon_{\mathrm{ff}}$ in the last several million years of their life cycles. Such a trend is consistent with the results in Grudić et al. (2019). Among these four model variations, GMCs with higher $M_{\mathrm{max,baryon}}$ exhibit higher $\varepsilon_{\mathrm{ff}}$. The high $\varepsilon_{\mathrm{ff}}$ (larger than $2\%$) in more massive GMCs can sustain for the entire life cycles of these GMCs, while less massive GMCs can have extremely low $\varepsilon_{\mathrm{ff}}$ (less than $2\%$) in the first few million years. Compared with both the SFE1 run and the Rad run, the star formation in the SN run is slightly more active across all mass ranges. Given that the selected timescale is $10$ Myr here, it is relatively short for the supernova to take effects thus the star formation in the SN run is mainly regulated by the galactic environment.
@@ 4.1 Comparisons to previous work
During the past several years, thanks to the increasing power of supercomputers and a better understanding of the ISM physics in galaxy formation simulations, a few studies have investigated the GMC properties from large-scale galaxy simulations. Here we would like to compare our results with them. Regarding the physical properties of identified GMCs, such as mass, virial parameter, effective radius, velocity dispersion, and Larson’s relations, GMCs identified in the Feedback ($3$D) case from Grisdale et al. (2018) in general exhibit similar distributions as shown in our fiducial run. The mass distribution of GMCs in the m12m case of Benincasa et al. (2020) is also similar to that in our fiducial run. Guszejnov et al. (2020) reports similar mass distribution of GMCs and similar Larson-like relation between $\sigma_{v}$ and $R_{\mathrm{eff}}$ in cosmological simulations.
@@
In terms of GMC lifetimes, Richings & Schaye (2016), Benincasa et al. (2020), J21, and our work all find that GMC lifetimes range from a couple of million years to several tens of million years. While the authors do not explicitly discuss trends between GMC lifetime and mass, in Richings & Schaye (2016), we note that their fiducial tracking scheme present a positive correlation between GMC lifetimes and masses, consistent with our conclusions. However, Benincasa et al. (2020) report limited dependence of the lifetime on the GMC mass and J21 proposes a negative correlation between GMC lifetimes and masses, which are contradictory to our findings on the positive correlation between GMC lifetimes and masses. In Benincasa et al. (2020), they used a relatively simple tracking algorithm and only followed strictly gravitationally bound GMCs. This approach might overlook the early stages of the GMC life cycles and miss key accretion phases, which are crucial to establishing the positive correlation between GMC lifetimes and masses. As for J21, though they introduced a similar tree network approach to capture both the temporal evolution of clouds and their interactions like our work, here we highlight one significant difference between J21’s and our GMC-tracking method, which may cause different results on the relation of GMC lifetimes and masses. Using the quasi-Lagrangian nature of arepo, we recorded additional information about the gas mass flow between nodes, which provided physical guidelines for selecting paths among multiple branches. J21, on the contrary, assumed the same probabilities for branches, regardless of the gas mass flowing through them. In this way, the GMC tracking in J21 might be misled by statistically overweighting minor branches of GMC trees. A massive GMC may go through many splitting or merging processes within its lifetime, and thus produce many minor branches. According to their tracking algorithm, these minor branches shorten the characteristic lifetimes of massive GMCs. It should be mentioned that in the GRIFFIN project, only clouds with the longest lifetime can form the most massive star clusters (Lahén et al. 2020), consistent with our findings here.
@@
Regarding the star formation process, while many studies focus on isolated GMC simulations to explore local SFE, our approach captures the influence of galactic environments self-consistently, while still enabling investigations of local SFE. In Section 3.5, we notice that the relation between the cloud surface density and the SFE is in a similar form as found in the isolated GMC simulations (e.g., Grudić et al. 2018; Li et al. 2019). Moreover, the large variations in the local SFE are observed in all simulation runs, consistent with Grisdale et al. (2019). Besides the fiducial run with the stellar feedback, we utilize the GMCs in the Nofeed run to quantitatively approximate the effect of the galactic shear on the local star formation efficiency with $\Sigma_{\mathrm{shear}}$. The best-fitting value of $\Sigma_{\mathrm{shear}}$ is comparable to the best-fitting value of $\Sigma_{\mathrm{fb}}$, suggesting the non-negligible role that the galactic shear plays in the star formation activity. We would like to point out that in Grudić et al. (2018) and Grudić et al. (2023), while their best-fitting $\Sigma_{\mathrm{fb}}$ is close to our best-fitting value, an extra parameter namely $\epsilon_{\mathrm{max}}$ is artificially introduced to account for any possible additional suppression of star formation other than stellar feedback. This approximation may partly mimic the effect of the $\Sigma_{\mathrm{shear}}$ in our study. Similar to our work, Khullar et al. (2024) also track life cycles of GMCs and study the GMC-scale SFE in galactic scale simulations, they get similar results on the importance of stellar feedback regulating star formation and also the large scatter in local SFE.
@@ 4.2 Limitations and future perspectives
We acknowledge several limitations of this work and outline potential directions for future research. First, while arepo, as a moving-mesh code, has a quasi-Lagrangian nature, it is not purely Lagrangian. The timescale for cell refinement and de-refinement is much shorter than the snapshot output cadence (i.e., $1\,\mathrm{Myr}$). Consequently, using cell IDs to track mass flow between GMCs across different epochs can lead to inaccuracies due to cell reconstruction during these intervals. To address this, future simulations will incorporate the Monte Carlo tracer particles (Genel et al. 2013) to allow the mass flow of clouds to be tracked more accurately. Second, though we find that galactic shear has a significant effect on GMC-scale star formation, it is yet unclear how such an effect would vary in different galactic environments. To explore how various galactic environments might have different effects on GMC-scale star formation, a companion study (Deng et al., in prep.) is currently underway to explore the evolution of GMCs in dwarf galaxies, utilizing the solar-mass-resolution ISM model, RIGEL (Deng et al. 2024).
@@ 5 Summary
In this study, we tracked the temporal evolution of a large sample of GMCs within simulated Milky Way-sized galaxies using the hydrodynamic code arepo under the SMUGGLE galaxy formation framework. The high mass resolution ($\sim 10^{3}M_{\odot}$) and high-cadence output frequency ($1$ Myr) enable the identification of cloud candidates and the construction of evolutionary trees. These evolutionary trees successfully capture the full GMC life cycle, including formation, gas accretion and collapse, star and cluster formation, and cloud dispersal. To investigate how different physical processes influence GMC evolution and star formation on cloud scales, we explore a set of model variations with different star formation and stellar feedback parameters. Below, we summarize the key findings:
@@
  • GMC lifetimes vary significantly across model variations, ranging from a few to several tens of million years, suggesting a strong influence from the star formation and stellar feedback subgrid models. Higher $\epsilon_{\rm ff}$ leads to faster gas consumption and cloud dispersal due to earlier and stronger stellar feedback and thus shorter lifetime. In the Nofeed run, GMCs live substantially longer than in other models, with over ten surviving beyond $20\,\mathrm{Myr}$. In the SN run, GMCs with lifetimes below $20\,\mathrm{Myr}$ tend to last longer than in the fiducial run, whereas those exceeding $20\,\mathrm{Myr}$ have shorter lifetimes. Conversely, in the Rad run, GMCs with lifetimes below $20\,\mathrm{Myr}$ show distributions similar to the fiducial run, while those above $20\,\mathrm{Myr}$ persist longer due to the absence of SN feedback.
@@
  • We capture the entire life cycles of GMCs in our simulated galaxies, from accretion and star formation to stellar feedback-driven cloud dispersal. We find that the accretion process continuously supplements the gas content throughout most of GMCs’ life cycles, making the cloud mass higher and virial parameter lower in the early stage. When the cloud mass reaches its peak, feedback from young stars disperses the cloud, reduces the cloud mass and inflates the virial parameter. Interestingly, there is a positive correlation between GMC lifetimes and GMC maximum gaseous masses, which stands true regardless of star formation and stellar feedback models.